# The Silent Majority: Demystifying Memorization Effect in the Presence of Spurious Correlations

### v2

@@ Abstract
Machine learning models often rely on simple spurious features – patterns in training data that correlate with targets but are not causally related to them, like image backgrounds in foreground classification. This reliance typically leads to imbalanced test performance across minority and majority groups. In this work, we take a closer look at the fundamental cause of such imbalanced performance through the lens of memorization, which refers to the ability to predict accurately on atypical examples (minority groups) in the training set but failing in achieving the same accuracy in the testing set. This paper systematically shows the ubiquitous existence of spurious features in a small set of neurons within the network, providing the first-ever evidence that memorization may contribute to imbalanced group performance. Through three experimental sources of converging empirical evidence, we find the property of a small subset of neurons or channels in memorizing minority group information. Inspired by these findings, we articulate the hypothesis: the imbalanced group performance is a byproduct of “noisy” spurious memorization confined to a small set of neurons. To further substantiate this hypothesis, we show that eliminating these unnecessary spurious memorization patterns via a novel framework during training can significantly affect the model performance on minority groups. Our experimental results across various architectures and benchmarks offer new insights on how neural networks encode core and spurious knowledge, laying the groundwork for future research in demystifying robustness to spurious correlation.
@@ 1 Introduction
Machine learning models often achieve high overall performance yet struggle in minority groups due to spurious correlations – patterns that align with the class label in training data but have no causal relationship with the target Sagawa et al. (2020a); Geirhos et al. (2020). For example, considering the task of distinguishing cows from camels in natural images, it is common to find 95% of cow images with grass backgrounds and 95% of camel images on sand. Models trained using standard Empirical Risk Minimization (ERM) often focus on minimizing average training error by depending on the spurious background attributes (“grass” or “sand”) instead of the core features (“cow” or “camel”). In such settings, models may yield good average accuracy but lead to high error rates minority groups (“cows on sand” or “camel on grass”) Ribeiro et al. (2016); Beery et al. (2018). This illustrates a fundamental issue: even well-trained models can develop systematic biases from these spurious attributes in their data, thus leading to alarmingly consistent performance drop for minority groups where the spurious correlation does not hold. Indeed, in Figure 1, we show the training and test accuracy on majority and minority groups of the Waterbirds benchmark for two popular models: ResNet-50 (He et al., 2016) and ViT-small (Dosovitskiy et al., 2021). We observe that majority groups have a smaller gap between the training and testing accuracy, as compared to minority groups which have a more significant gap. Thus, understanding the underlying causes of such imbalanced performance between majority and minority groups is crucial for their reliable and safe deployment in various real-world scenarios Blodgett et al. (2016); Buolamwini and Gebru (2018); Hashimoto et al. (2018).
@@
Figure 1: Imbalanced Group Performance on Waterbirds. Majority groups ($\mathcal{G}_{0}$ and $\mathcal{G}_{3}$) show a minimal gap between training and test accuracy, while minority groups ($\mathcal{G}_{1}$ and $\mathcal{G}_{2}$) yield a significantly larger discrepancy. Both models are trained with ERM.
@@
The minority groups are atypical examples to neural networks (NNs), as these small subsets of examples bear a resemblance to majority groups due to the same spurious attribute but have distinct labels. Recent efforts have shown that NNs often ‘memorize’ atypical examples, primarily in the final few layers of the model Baldock et al. (2021); Stephenson et al. (2021), and possibly even in specific locations of the model Maini et al. (2023). Memorization, in this context, is defined as the neural network’s ability to accurately predict outcomes for atypical examples (e.g., mislabeled examples) in the training set through ERM training. This is in striking analogy to the spurious correlation issue, because 1) the minority examples are atypical examples by definition, and 2) the minority examples are often more accurately predicted during training but poorly predicted during testing (as demonstrated in Figure 1). Therefore, a natural open question arises: Does memorization play a role in spurious correlations?
@@
To achieve these goals, we show the existence of a phenomenon named spurious memorization. We define ‘spurious memorization’ as the ability of NNs to accurately predict outcomes for atypical (i.e., minority) examples during training by deliberately memorizing them in certain part of the model. Indeed, we first identify that a small set of neurons is critical for memorizing minority examples. These critical neurons significantly affect the model performance on minority examples during the training, but only have minimal influence on majority examples. Furthermore, we show that these critical neurons only account for a very small portion of the model parameters. Such a memorization by a small portion of neurons causes the model performance on minority examples to be non-robust, which leads to the poor testing accuracy on minority examples despite the high training accuracy. Overall, our study offers a potential explanation for the differing performance patterns of NNs when handling majority and minority examples.
Figure 2: Group accuracy change by pruning Top-$k$ neuron(s) with gradient-based (left) and magnitude-based (right) criterion. Within each group $\mathcal{G}$, three bars with gradated hues indicate the accuracy shift after zeroing out the neurons with the top-1, top-2, and top-3 largest gradients or magnitudes, respectively.
Our systematic study is performed in two stages. In Stage I, to verify the existence of critical neurons, we identify two experimental sources to trace spurious memorization at the neuron and layer level. These two sources are unstructured tracing (assessing the role of neurons within the entire model for spurious memorization using heuristics including weight magnitude and gradient) and structured tracing (assessing the role of neurons within each individual layer with similar heuristics). Specifically, by evaluating the impact of spurious memorization via unstructured and structured tracing at the magnitude and gradient level (Section 3.1), we observe a substantial decrease in minority group accuracy, contrasting with a minimal effect on the majority group accuracy. This suggests that at unstructured and structured level, the learning of minority group opposes the learning of majority group, and indicates that 1) critical neurons for spurious memorization indeed exist within NNs; 2) both gradient and magnitude criteria are effective tools for identifying these critical neurons; and 3) NNs tend to memorize typical examples from majority groups on a global scale, whereas the memorization of minority examples is localized to a miniature set of nodes (i.e. critical neurons). Overall, we provide converging empirical evidence to confirm the existence of critical neurons for spurious memorization.
@@
2 Related Work
Spurious Correlation.
Machine learning systems often exhibit imbalanced performance across different groups in the presence of spurious correlations, where minority groups tend to have significantly lower test accuracy than majority groups, as contrast to the training phase where the training accuracy of both groups tend to be more evenly balanced Sagawa et al. (2020a); Geirhos et al. (2020). Recent works Geirhos et al. (2020); Sagawa et al. (2020b); Shah et al. (2020); Hermann and Lampinen (2020) on this area reveal that NNs suffer from extreme simplicity bias and tend to prioritize simple spurious features, while ignoring the core features.
A large body of previous works aim to enhance robustness against spurious correlations. On one hand, when group labels are available, researchers have developed methods to balance groups of different sizes during training, employing strategies like class balancing He and Garcia (2009); Cui et al. (2019), importance weighting Shimodaira (2000); Byrd and Lipton (2019), robust optimization Sagawa et al. (2020a); Kirichenko et al. (2023); Izmailov et al. (2022), or contrastive learning Taghanaki et al. (2021). On the other hand, when group labels are unavailable, a common theme is to initially train an auxiliary model using ERM, and then utilizing the predictions of the auxiliary model to infer group information. Subsequently, they often train a second, more robust model, employing strategies such as sample balancing Liu et al. (2021); Nam et al. (2020), robust optimization Creager et al. (2021); Nam et al. (2022); Sohoni et al. (2020), or contrastive learning Zhang et al. (2022); Zhang and Ré (2022); Yang et al. (2023); You et al. (2024) with the inferred group labels.
In this work, we pioneer a thorough investigation to reveal the mechanisms that cause spurious correlations. Despite effectively boosting the worst-group accuracy, there still exists a large gap in the reliability of the existing methods due to a lack of understanding of the underlying mechanisms causing spurious correlations. Existing studies show that spurious correlations can be effectively tackled by just retraining the model’s last layer with a small subset of data where group sizes are balanced Kirichenko et al. (2023); Izmailov et al. (2022). This inspires an intriguing question: Could the imbalanced group performance observed between typical majority examples and atypical minority examples be attributed to certain yet-to-be-uncovered memorization mechanisms within NNs?
Memorization.
Prior work aims at explaining the interplay between memorization and robustness in the context of the long-tailed nature of datasets. Recent studies Feldman and Zhang (2020); Feldman (2020) show that certain examples are repeatedly forgotten during training, while others are easily memorized. Such forgettable examples often stem from semantically ambiguous or underrepresented groups with rare features in the training set. Echoing the “early learning phenomenon”, recent finding Frankle et al. (2020) suggests that simpler examples are learned fast in early training stages. Based on this understanding, some works utilize learning time as a metric to determine an example’s memorization status (Carlini et al., 2019; Jiang et al., 2020). Furthermore, other works focus on the concept of forgetting time, indicating that examples frequently forgotten during training Toneva et al. (2018) or quickly forgotten when held out are those that are memorized Maini et al. (2022).
There is a growing body of research investigating the location of memorization in NNs (Baldock et al., 2021; Stephenson et al., 2021). Further, there has been growing interest in how modifications to a miniature set of neurons within NNs can alter predictions, particularly in understanding where atypical example information is stored in the network Meng et al. (2022); Zhu et al. (2020). In addition, recent work Sinitsin et al. (2020) developed methods to rectify neural network errors by adjusting a few neurons. Very recent work Maini et al. (2023) also investigated the minimum number of neurons required to change a network’s prediction. In contrast, we focus on study fundamental properties of memorization in the presence of spurious correlation. A different definition of memorization has also been studied (Sagawa et al., 2020b), where it is defined as a situation where an algorithm memorizes the label $y_{i}$ if its prediction for $x_{i}$, based on the rest of the dataset, changes significantly once the pair $(x_{i},y_{i})$ is included in the training dataset.
In summary, our work is fundamentally different from prior works by highlighting two key insights: 1) a miniature set of neurons plays a pivotal role in memorizing spurious correlation; 2) simply eliminating these critical neurons significantly improves group robustness. To the best of our knowledge, this work is the first to systematically investigate the role of critical neurons in causing imbalanced group performance and analyzing the mechanism of spurious memorization within these neurons. Further, we propose to redirect the influence of these critical neurons to the entire network when spurious memorization occurs, thereby mitigating imbalanced performance across different groups.
@@ 3.1.1 Unstructured Tracing
Figure 3: Largest gradient neurons’ magnitude ranking distribution (left) and largest magnitude neurons’ gradient ranking distribution (right). The x-axis percentage refers to the relative ranking of neurons’ gradients or magnitudes across the entire network (0%: smallest, 100%: largest). Full details in Section 5.1.
To assess the importance of these selected neurons in memorizing examples, we zero them out and calculate the change in group accuracy on the training set. The pruned model is identified as $f(\mathbf{m}_{j}\odot\bm{\theta},\cdot)$, where $\mathbf{m}_{j}$ is a mask with neurons in $\mathcal{I}_{j}$ being masked. The change in accuracy $\Delta_{\textnormal{acc}}$ for each group $j$ is given by $\Delta_{\textnormal{acc}}(j)=|\textnormal{acc}(\mathcal{D}_{j},f(\bm{\theta},\cdot))-\textnormal{acc}(\mathcal{D}_{j},f(\mathbf{m}_{j}\odot\bm{\theta},\cdot))|$, where acc represents the accuracy. In our experiments below, all the group accuracy change is based on 97.34% ($\mathcal{G}_{0}$), 47.83% ($\mathcal{G}_{1}$), 69.64% ($\mathcal{G}_{2}$), 97.63% ($\mathcal{G}_{3}$).
@@
In Figure 4 and  5, we found that 1) the results from random initialization ($\mathbf{\sigma}=0.005$) closely resemble those from the pruning method. Notably, the minority groups show a $2\%-12\%$ shift in group accuracy compared to the majority groups’ $0\%-2.6\%$ shift, since random initialization converges to pruning as the standard deviation of the Gaussian random variables $\{\epsilon_{i}\}_{i\in\mathcal{I}_{j}}$ decreases; 2) with random noise added ($\mathbf{\sigma}=0.005$), the accuracy changes in minority groups still surpass those in majority groups. We also observe that the extent of accuracy change with random noise is much smaller than that observed with random initialization and pruning. This occurs because, in the presence of random noise, neuron values are not reset to a mean-zero state, allowing their initial values to impact the model’s performance across different groups. Additional experimental results on random initialization and random noise with various standard deviations can be found below.
@@
Overall, regardless of the scale variation in the accuracy shifts, our experiments using pruning, random initialization, and random noise consistently demonstrate that the accuracy of minority groups is significantly sensitive to the alteration of a handful of selected neurons. This finding confirms the existence of a small subset of critical neurons for memorizing minority examples during training. Moreover, it validates that both gradient-based and magnitude-based criteria are effective in identifying these critical neurons.
@@ 3.2 Spurious Memorization by Critical Neurons
In Section 3.1, our experiments have established the presence of a small set of critical neurons responsible for memorizing minority examples during training. This underscores the role of spurious memorization as a significant factor in imbalanced group performance. In this section, we take a further step in demystifying the cause of imbalanced group performance under spurious correlation, particularly focusing on the discrepancy in the test accuracy between majority and minority groups.
To further validate the hypothesis that spurious memorization is a key factor in the imbalanced group performance, we investigate whether countering spurious memorization during training could lead to improved test accuracy on minority groups. Our findings affirmatively answer this question. By specifically targeting and removing spurious memorization via a specialized fine-tuning framework, we observe a consistent improvement in the test accuracy for minority groups. We report extensive experimental results across different model architectures, including ResNet-50 and ViT-Small, and on benchmark datasets including Waterbirds (Sagawa et al., 2020a; Wah et al., 2011) and CelebA (Liu et al., 2015), providing comprehensive analysis on the effects of spurious memorization on imbalanced group performance.
@@ 3.2.1 Interference with Spurious Memorization
Consider an arbitrary input image $\mathbf{x}$, and denote by $\mathbf{x}^{\prime}$ and $\mathbf{x}^{\prime\prime}$ two augmentations of $\mathbf{x}$. The loss term on the input $\mathbf{x}$ (together with its positive and negative pairs; see Section 5.1 for details) can be formulated as:

### v3

# Uncovering Memorization Effect in the Presence of Spurious Correlations

@@ Abstract
Machine learning models often rely on simple spurious features – patterns in training data that correlate with targets but are not causally related to them, like image backgrounds in foreground classification. This reliance typically leads to imbalanced test performance across minority and majority groups. In this work, we take a closer look at the fundamental cause of such imbalanced performance through the lens of memorization, which refers to the ability to predict accurately on atypical examples (minority groups) in the training set but failing in achieving the same accuracy in the testing set. This paper systematically shows the ubiquitous existence of spurious features in a small set of neurons within the network, providing the first-ever evidence that memorization may contribute to imbalanced group performance. Through three experimental sources of converging empirical evidence, we find the property of a small subset of neurons or channels in memorizing minority group information. Inspired by these findings, we hypothesize that spurious memorization, concentrated within a small subset of neurons, plays a key role in driving imbalanced group performance. To further substantiate this hypothesis, we show that eliminating these unnecessary spurious memorization patterns via a novel framework during training can significantly affect the model performance on minority groups. Our experimental results across various architectures and benchmarks offer new insights on how neural networks encode core and spurious knowledge, laying the groundwork for future research in demystifying robustness to spurious correlation. Our codes are available in here.
@@ 1 Introduction
Machine learning models often achieve high overall performance, yet struggle in minority groups due to spurious correlations – patterns that align with the class label in training data but have no causal relationship with the target Sagawa et al. (2020); Geirhos et al. (2020). For example, considering the task of distinguishing cows from camels in natural images, it is common to find 95% cow images with grass backgrounds and 95% of camel images on sand. Models trained using standard Empirical Risk Minimization (ERM) often focus on minimizing the average training error by depending on spurious background attributes (“grass” or “sand”) instead of the core characteristics (“cow” or “camel”). In such settings, models may yield good average accuracy but lead to high error rates in minority groups (“cows on sand” or “camel on grass”) Ribeiro et al. (2016); Beery et al. (2018). This illustrates a fundamental issue: even well-trained models can develop systematic biases from these spurious attributes in their data, thus leading to alarmingly consistent performance drop for minority groups where the spurious correlation does not hold. Indeed, in Figure 1, we present both the training and test accuracy on the majority and minority groups of the Waterbirds benchmark for two popular models: ResNet-50 (He et al., 2016) and ViT-small (Dosovitskiy et al., 2021). It is clear from Figure 1 that the test performance is poor on minority groups (1 and 2). Moreover, we observe that majority groups have a smaller gap between the training and testing accuracy, as compared to minority groups that have a more significant gap. Thus, understanding the underlying causes of this unbalanced performance between the majority and minority groups is crucial to their reliable and safe deployment in various real-world scenarios Blodgett et al. (2016); Buolamwini and Gebru (2018); Hashimoto et al. (2018).
@@
Figure 1: Imbalanced Group Performance on Waterbirds. Majority groups ($\mathcal{G}_{0}$ and $\mathcal{G}_{3}$) show the minimal gap between training and test accuracy, while minority groups ($\mathcal{G}_{1}$ and $\mathcal{G}_{2}$) yield a significantly larger discrepancy. Both models are trained with ERM. Here the star superscript (*) in the figure is used to emphasize the minority groups $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
@@
The minority groups are atypical examples to neural networks (NNs), as these small subsets of examples bear a similarity to majority groups due to the same spurious attribute, but have distinct labels. Recent efforts have shown that NNs often ‘memorize’ atypical examples, primarily in the final few layers of the model Baldock et al. (2021); Stephenson et al. (2021), and possibly even in specific locations of the model Maini et al. (2023). Memorization, in this context, is defined as the neural network’s ability to accurately predict outcomes for atypical examples (e.g., mislabeled examples) in the training set through ERM training. This is in striking analogy to the spurious correlation issue, because 1) the minority examples are atypical examples by definition, and 2) the minority examples are often more accurately predicted during training but poorly predicted during testing (as demonstrated in Figure 1). Therefore, a natural open question arises: Does memorization play a role in spurious correlations?
@@
To achieve these goals, we show the existence of a phenomenon named spurious memorization. We define ‘spurious memorization’ as the ability of NNs to accurately predict outcomes for atypical (i.e., minority) examples during training by deliberately memorizing them in certain part of the model. Indeed, we first identify that a small set of neurons is critical for memorizing minority examples. These critical neurons significantly affect the model performance on minority examples during training, but only have minimal influence on majority examples. Furthermore, we show that these critical neurons only account for a very small portion of the model parameters. Such a memorization by a small portion of neurons causes the model performance on minority examples to be non-robust, which leads to the poor testing accuracy on minority examples despite the high training accuracy. Overall, our study offers a potential explanation for the differing performance patterns of NNs when handling majority and minority examples.
Figure 2: Group accuracy change by pruning Top-$k$ neuron(s) with gradient-based (left) and magnitude-based (right) criterion. Within each group $\mathcal{G}$, three bars with gradated hues indicate the accuracy shift after zeroing out the neurons with the top-1, top-2, and top-3 largest gradients or magnitudes, respectively. Note that the minority groups $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are emphasized with the star superscript (*).
Our systematic study is performed in two stages. In Stage I, to verify the existence of critical neurons, we identify two experimental sources to trace spurious memorization at the neuron and layer level. These two sources are unstructured tracing (assessing the role of neurons within the entire model for spurious memorization using heuristics including weight magnitude and gradient) and structured tracing (assessing the role of neurons within each individual layer with similar heuristics). Specifically, by evaluating the impact of spurious memorization via unstructured and structured tracing at the magnitude and gradient level (Section 2.1), we observe a substantial decrease in minority group accuracy, contrasting with a minimal effect on the majority group accuracy. This suggests that at unstructured and structured level, the learning of minority group opposes the learning of majority group, and indicates that 1) critical neurons for spurious memorization indeed exist within NNs; 2) both gradient and magnitude criteria are effective tools for identifying these critical neurons; and 3) NNs tend to memorize typical examples from majority groups on a global scale, whereas a miniature set of nodes (i.e. critical neurons) is involved in the memorization of minority examples to a greater extent than other neurons. Overall, we provide converging empirical evidence to confirm the existence of critical neurons for spurious memorization.
@@ 2.1.1 Unstructured Tracing
Figure 3: Largest gradient neurons’ magnitude ranking distribution (left) and largest magnitude neurons’ gradient ranking distribution (right). The x-axis percentage refers to the relative ranking of neurons’ gradients or magnitudes across the entire network (0%: smallest, 100%: largest).
To assess the importance of these selected neurons in memorizing examples, we zero them out and calculate the change in group accuracy on the training set. The pruned model is identified as $f(\mathbf{m}_{j}\odot\bm{\theta},\cdot)$, where $\mathbf{m}_{j}$ is a mask with neurons in $\mathcal{I}_{j}$ being masked. The change in accuracy $\Delta_{\textnormal{acc}}$ for each group $j$ is given by $\Delta_{\textnormal{acc}}(j)=|\textnormal{acc}(\mathcal{D}_{j},f(\bm{\theta},\cdot))-\textnormal{acc}(\mathcal{D}_{j},f(\mathbf{m}_{j}\odot\bm{\theta},\cdot))|$, where acc represents the accuracy. In our experiments below, all the group accuracy change is based on the following training accuracy: 97.34% ($\mathcal{G}_{0}$), 47.83% ($\mathcal{G}_{1}$), 69.64% ($\mathcal{G}_{2}$), 97.63% ($\mathcal{G}_{3}$)  [^2] For completeness, we also report the baseline test accuracy: 96.98% ($\mathcal{G}_{0}$), 35.68% ($\mathcal{G}_{1}$), 56.98% ($\mathcal{G}_{2}$), and 96.26% ($\mathcal{G}_{3}$). The baseline test accuracy follows the same pattern as the training accuracy..
@@
Refer to caption Refer to caption
In Figure 4 and  5, we found that 1) the results from random initialization ($\mathbf{\sigma}=0.005$) closely resemble those from the pruning method. Notably, the minority groups show a $2\%-12\%$ shift in group accuracy compared to the majority groups’ $0\%-2.6\%$ shift, since random initialization converges to pruning as the standard deviation of the Gaussian random variables $\{\epsilon_{i}\}_{i\in\mathcal{I}_{j}}$ decreases; 2) with random noise added ($\mathbf{\sigma}=0.005$), the accuracy changes in minority groups still surpass those in majority groups. We also observe that the extent of accuracy change with random noise is much smaller than that observed with random initialization and pruning. This occurs because, in the presence of random noise, neuron values are not reset to a mean-zero state, allowing their initial values to impact the model’s performance across different groups. Moreover, we conduct the additional unstructured tracing experiment on CelebA, as shown in Table 11 (in Supplementary materials). We can see that the results for the minority group of CelebA align with those obtained on the Waterbirds dataset. As shown in Table 13, our results in the Waterbirds dataset reveal that when modifying critical neurons, the training accuracy for minority groups ($\mathcal{G}_{1}$ and $\mathcal{G}_{2}$) consistently drops across nearly all experimental setups, whereas the corresponding test accuracy remains relatively stable. This discrepancy reinforces the interpretation that these neurons play a memorization role. Furthermore, the raw values reported in Table 14 further substantiate these findings, showing that for minority groups, training accuracy decreases consistently in all experimental settings and for every choice of $k$ in the top-$k$ analysis. Together, these results provide compelling evidence that identified critical neurons are primarily responsible for memorization in minority groups, rather than affecting overall generalization. Additional experimental results on random initialization and random noise with various standard deviations can be found below.
Refer to caption Refer to caption
@@
Overall, regardless of the scale variation in the accuracy shifts, our experiments using pruning, random initialization, and random noise consistently demonstrate that the accuracy of minority groups is significantly sensitive to the alteration of a handful of selected neurons. This finding suggests that a small subset of critical neurons contributes more significantly to the memorization of minority examples during training than other neurons. Moreover, it validates that both gradient-based and magnitude-based criteria are effective in identifying these critical neurons.
@@ 2.2 Spurious Memorization by Critical Neurons
In Section 2.1, our experiments have empirically demonstrated the presence of a small set of critical neurons involved in the memorization of minority examples during training. This underscores the role of spurious memorization as a significant factor in imbalanced group performance. In this section, we take a further step in demystifying the cause of imbalanced group performance under spurious correlation, particularly focusing on the discrepancy in the test accuracy between majority and minority groups.
To further validate the hypothesis that spurious memorization is a key factor in the imbalanced group performance, we investigate whether countering spurious memorization during training could lead to improved test accuracy on minority groups. Our findings affirmatively answer this question. By specifically targeting and removing spurious memorization via a specialized fine-tuning framework, we observe a consistent improvement in the test accuracy for minority groups. We report extensive experimental results across different model architectures, including ResNet-50 and ViT-Small, and on benchmark datasets including Waterbirds (Sagawa et al., 2020; Wah et al., 2011) and CelebA (Liu et al., 2015), providing comprehensive analysis on the effects of spurious memorization on imbalanced group performance.
@@ 2.2.1 Interference with Spurious Memorization
Consider an arbitrary input image $\mathbf{x}$, and denote by $\mathbf{x}^{\prime}$ and $\mathbf{x}^{\prime\prime}$ two augmentations of $\mathbf{x}$. The loss term on the input $\mathbf{x}$ (together with its positive and negative pairs) can be formulated as: